# Cryptocurrency as an Investable Asset Class: Coming of Age

### v2

@@ Abstract
Cryptocurrencies are coming of age. We organize empirical regularities into ten stylized facts and analyze cryptocurrency through the lens of empirical asset pricing. We find important similarities with traditional markets—risk-adjusted performance is broadly comparable, and the cross-section of returns can be summarized by a small set of factors. However, cryptocurrency also has its own distinct character: jumps are frequent and large, and blockchain information helps drive prices. This common set of facts provides evidence that cryptocurrency is emerging as an investable asset class.
@@ 1 INTRODUCTION
This survey establishes ten key empirical regularities of cryptocurrencies motivated by the view that cryptocurrency is an asset and thus can and should be analyzed from the asset-pricing point of view and, more specifically, from the cryptocurrency asset pricing view. Although the literature is relatively young—cryptocurrencies trace back to Nakamoto (2008)’s short paper on Bitcoin—it is now at the point of maturity where it is useful to establish a core of stylized facts in cryptocurrency markets that can be used for both empirical and theoretical work. This common set of facts also provides emerging evidence that cryptocurrency is becoming an investable asset class.
@@
When the paper of Nakamoto (2008) introduced a permissionless blockchain that provides public record-keeping without a central authority, financed and governed by a native token, the canonical literature on money had reached a degree of maturity. Permissionless ledgers can be viewed as technologies that supply public memory, with token systems financing and governing that record—keeping, and providing histories, communication, and enforcement mechanisms that are required for credit. With the rise of cryptocurrencies, these new ideas began to emerge primarily from practice more than theory and reopened classic monetary questions—record-keeping, commitment, enforcement—but now viewed from the point of view of blockchain economics.
@@
When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory. The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing. A sizable literature uses an asset pricing lens to study risk exposures, cross-sectional return patterns, cash-flow-like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020, Liu et al. 2022, Alvarez et al. 2023, Sockin & Xiong 2023a, von Luckner et al. 2023, Kogan et al. 2024).
@@
Many have viewed cryptocurrencies—and, not unjustifiably, many still view them—as hype, speculation, or fraud. When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system. [^1] Foley et al. (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations. It is worth noting that several important financial innovations, such as the limited-liability company and mortgage-backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
@@
Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018, Leshno & Strack 2020, Biais et al. 2019, Kogan et al. 2022, Brunnermeier et al. 2019, Benigno et al. 2022, Capponi et al. 2025, Jermann & Xiang 2025), regulation (Cong & He 2019,Townsend 2020, Auer et al. 2025), digital asset pricing and valuation (Athey et al. 2016, Cong et al. 2021b, Biais et al. 2023), transaction costs (Easley et al. 2019, Cong et al. 2021a, Basu et al. 2023), central bank money (Auer et al. 2022, Whited et al. 2023, Schilling et al. 2024), securitization (Cong et al. 2022, Sockin & Xiong 2023a, Rogoff & You 2023, Goldstein et al. 2024), safe assets (Ma et al. 2025), international capital flow (von Luckner et al. 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al. 2024).
@@
We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of ten stylized empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes. These empirical facts are also useful to evaluate and revisit both the classic and the new theories of money.
@@
Ten Facts in Cryptocurrency Asset Pricing and Empirical Finance. Facts 1 and 2 overview the core asset-pricing characteristics of cryptocurrency markets—the risk-return trade-off and the diversification benefits. Facts 3 and 4 discuss the structure of returns in cryptocurrency—factors, or “smart-beta”, which explain the cross-section of returns and optimal portfolio shares. Fact 5 shows the importance of large, discrete price movements—jumps—and the importance of extreme events for understanding cryptocurrency risk and valuation. Fact 6 shows that the higher-order terms of a parsimonious four-factor model account for a large fraction of the cross-section of returns and provide a “glass box” alternative to the “black box” neural net models with interpretable economic factor structure. Fact 7 shows the importance of the blockchain structure and network economy phenomena for cryptocurrency prices—on-chain user adoption influences cryptocurrency returns. Facts 8 documents large and persistent market inefficiencies due to frictions that limit arbitrage, and discuss risks for arbitrageurs. Fact 9 is about the cryptocurrency futures market and funding premia, and their evolution. Finally, Fact 10 is about the importance of regulation, disclosure, and transparency. In sum, these ten facts can be viewed as providing evidence that cryptocurrency, and more broadly, digital assets are coming of age as an emerging investable asset class.
@@ 2 DATA
Second, for the construction of the crypto carry strategy of Schmeling et al. (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025. The data on new addresses are from Liu et al. (2021). Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025. Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.222Finally, we obtain price indices for the U.S. stock market, U.S. corporate and government bonds, a broad commodity index, gold, the U.S. dollar index, the Dow Jones US Real Estate Index, the U.S. constant maturity 10-year nominal and real bond yields from Bloomberg. The corresponding Bloomberg tickers are SPX Index, LUACTRUU Index, LUATTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
@@
Unless otherwise noted, we aggregate daily closing prices to construct weekly coin returns. Each year is divided into 52 weeks: the first week contains the year’s first seven days, the next 50 weeks consist of seven days each, and the final week contains the remaining days of the year. For the construction of weekly returns, we exclude coins with a market capitalization below \$1 million at the formation date and winsorize returns at the 0.0025% level in both tails. We construct a cryptocurrency market return as the value-weighted return of all underlying coins. The cryptocurrency excess market return is constructed as the difference between the cryptocurrency market return and the risk-free rate measured as the 1-month T-bill rate at the weekly frequency, which we obtain from the Federal Reserve Bank of St. Louis.
@@
Figure 1 summarizes the time-series evolution of cryptocurrency prices and market capitalization. A one-dollar investment at the start of the sample grows to about \$160 in Bitcoin and \$80 in the overall market. There is substantial variation over time. Market capitalization reaches nearly \$4 trillion, with Bitcoin and Ethereum accounting for approximately 58% and 12% of the total at the end of the sample.
@@
3 THE TEN FACTS OF CRYPTOCURRENCY
We organize the main discussions of this review into the ten key stylized empirical facts on cryptocurrency markets.
3.1 Fact 1: High return, high volatility—normal Sharpe ratio
Fact 1: The cryptocurrency market has an order of magnitude higher returns and volatility than equities, but a broadly similar Sharpe ratio.
@@ 3.1 Fact 1: High return, high volatility—normal Sharpe ratio
Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable. Fact 1 revisits and confirms this finding in a longer and larger dataset up to September 2025.
Table 1, Panel A, reports summary statistics for the sample beginning in 2014 and ending in September 2025, as well as for the samples pre-2020 and post-2020. We compare returns on the aggregate coin market, a large-cap index of the top 100 cryptocurrencies (Coin 100), Bitcoin, and the aggregate U.S. stock market. The average returns and volatilities of all cryptocurrency return series are an order of magnitude greater than those of the stock market across all sample periods. For instance, average weekly returns exceed 1% for all cryptocurrency indices, compared with less than 0.3% for the aggregate stock market. However, the weekly Sharpe ratios are similar across cryptocurrencies and equities. This, confirms that the Sharpe ratio calculated by Liu & Tsyvinski (2021) holds in the most recent sample.
Table 1, Panel A, reports maximum drawdowns and run-ups for the four return indices. Drawdowns are important as they provide a measure of tail-risk not fully reflected by standard deviation. Drawdowns and run-ups are substantially larger for cryptocurrencies than for equities. Importantly, for all three cryptocurrency indices, maximum drawdowns and run-ups are smaller in magnitude in the post-2020 sample than in the pre-2020 sample, while the reverse holds for equities. This pattern can also be observed in the standard deviations of returns—they decline for cryptocurrencies and rise for equities in the more recent period.
@@
Notes: This table summarizes the main properties of coin market return. Panel A reports weekly summary statistics of the coin market index and compares them with the value-weighted index of the top 100 coins by market capitalization (Coin 100), Bitcoin and U.S. stock market returns. Reported measures include the mean, standard deviation, $t$-statistic, Sharpe ratio, skewness, kurtosis, maximum drawdown and run-up. Panel B presents pairwise correlation coefficients between coin market returns and returns on the U.S. stock market, corporate bonds, government bonds, a commodity index, gold, the dollar index, and the changes in inflation expectations. $t$-statistics are reported in parentheses. The full sample covers the period from January 7, 2014 to September 6, 2025. The pre-2020 sample is the sample before Jan 1, 2020, and the post-2020 sample is the sample since Jan 1, 2020.
3.2 Fact 2: Cryptocurrency is a distinct asset: it moves largely on its own—but increasingly with equities, gold and commodities
Fact 2: The aggregate cryptocurrency market remains largely uncorrelated with most traditional asset classes, but its correlation with U.S. equities has increased sharply—from 2% before 2020 to 37% after 2020.
How cryptocurrencies co-move with traditional asset classes is a first-order question for asset pricing and portfolio choice (e.g., Markowitz 1952). Correlations are an important determinant of diversification benefits and show whether cryptocurrency risk is distinct to that market or simply represent familiar risks from other markets. We revisit the cryptocurrency aggregate correlation structure using updated data, contrasting the pre-2020 and post-2020 return correlation.
One of the main results of Liu & Tsyvinski (2021) is that aggregate cryptocurrency returns have insignificant exposures to the returns of traditional asset classes, such as equities, bonds, commodities, and foreign exchanges. In other words, the correlation between the aggregate cryptocurrency returns and those of traditional asset classes are low. Fact 2 revisits this finding. We show that there is a rising correlation with the U.S. stock market.
Table 1, Panel B, reports correlations of aggregate cryptocurrency returns with those of major asset classes. In the pre-2020 sample, correlations with equities, corporate bonds, government bonds, commodities, gold, and the dollar are all statistically insignificant, confirming the earlier findings of Liu & Tsyvinski (2021). In contrast, in the post-2020 sample, aggregate cryptocurrency returns become significantly correlated with equities (37%), commodities (27%), gold (18%), and corporate bonds (13%) at the 5% level. The correlation with the dollar index is negative (-10%) and significant at the 10% level. Government bonds remain the only asset class with no significant link.
The most striking change is the correlation with the U.S. stock market, which rises from only 2% before 2020 to 37% afterwards. This substantial increase is an important indicator that cryptocurrencies have become more integrated with equity markets. We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S. Treasury bonds with 10-year maturity. Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years. In our view, this link with expected inflation is important, as it provides empirical evidence on the original motivation of cryptocurrencies as assets designed to hedge against the inflationary risks of government printing money.
The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong et al. (2023) and Jiang et al. (2025). Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization. The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market.
3.3 Fact 3: Significant diversification benefits from adding small amount of cryptocurrencies in investment portfolios
Fact 3: Cryptocurrency idiosyncratic risk can be significantly reduced by constructing optimal mean-variance asset portfolios that include cryptocurrencies—the inclusion of cryptocurrencies enlarges the mean-variance frontier. Optimal portfolio allocates 3.1% cryptocurrency for retail investors and 5.5% for institutional investors.
Modern finance theory, starting from Markowitz (1952), emphasizes the central insight of diversification: an asset’s covariance with the investor’s portfolio matters more than its volatility. Low correlations push the efficient frontier outward, and volatile assets can add value if their expected returns compensate for marginal risk. Since Fact 1 establishes that cryptocurrencies have low correlations with stocks and bonds, they can be effective risk diversifiers under standard portfolio-theory frameworks.
Liu & Tsyvinski (2021) show that cryptocurrencies can improve portfolio efficiency because of their historically low correlations with traditional asset classes and comparatively high expected returns. Using the Black–Litterman portfolio allocation model (Black & Litterman 1990) to incorporate investor views, Liu & Tsyvinski (2021) estimate that a representative investor would optimally allocate between 1% and 6% of wealth to cryptocurrencies, depending on the strength of those views. A complementary view comes from Bayesian portfolio theory. Duchin et al. (2025) obtain similar optimal weights in the 1%–5% range. Ang et al. (2023) find much larger optimal weights when investors have a preference for positive skewness, such as under constant relative risk aversion.
Because the optimal crypto weight is sensitive to the joint return distribution—especially the correlation between cryptocurrencies and traditional assets—and because this correlation structure shifted after 2020 (see Fact 2), we re-estimate the theoretical optimal allocations using post-2020 data under the same Black–Litterman framework.
Moreover, we extend the analysis to two representative investor classes: a retail investor and an institutional investor. In calculating the theoretical optimal crypto portfolio shares for them, we differ only in their investable options. This simplification abstracts from practical complications, such as liquidity, custody, and regulatory constraints, to make the exercise tractable.
For a retail investor, we start with a baseline portfolio that is 50–50 between financial assets and real estate, with the financial portion 60–40 stocks–bonds. Based on the model, if we add an allocation to the cryptocurrency market, we would raise the portfolio’s Sharpe ratio from 0.39 to 0.52, with an optimal crypto weight of 3.1% of total wealth. As a historical counterfactual, we also compute the theoretical optimal portfolio using only pre-2020 data; the optimal crypto share is 1.4%. Mapping this to per-capita U.S. net wealth in 2019, a 1.4% allocation implies an investment of about \$3,500; held from 2020 to the end of our sample, this position would have grown by roughly \$52,000 (about 1,462%). This is a back-of-the-envelope, buy-and-hold calculation that assumes no rebalancing.
For an institutional investor, we consider a broader investable universe—equities, government and corporate bonds, commodities, and real estate—with baseline weights proportional to each asset class’s market capitalization. Solving for the optimal portfolio and assuming slightly greater confidence in views on cryptocurrencies than in the retail case, the model implies an increase in the portfolio’s Sharpe ratio from 0.69 to 0.95, with an optimal share of 5.5% in the cryptocurrency market.
These results show that—even with the post-2020 increase in comovement with equities—optimal allocations to cryptocurrency are sizable and in the 3–5% range, with larger weights for institutional portfolios. It is important to note that the estimations in this section are back-of-the-envelope and rely on the properties of returns that may not persist and abstract from important issues related to liquidity, custody, and regulatory risk, and therefore should be viewed as theoretical upper bounds on the potential diversification benefits of this emerging investable asset class, not as actual investment recommendations. For context, CalPERS, one of the largest U.S. pension funds, reports portfolio shares of comparable magnitude in emerging-market debt (5.3%), high-yield bonds (5.3%), private debt (3.8%), and mortgage-backed securities (5.8%) as of June 2025, indicating that the model-implied weights of cryptocurrencies are within a realistic range.
3.4 Fact 4: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value

### v3

@@ Abstract
We organize existing empirical regularities of cryptocurrencies into seven stylized facts and analyze cryptocurrencies through the lens of empirical asset pricing. We find important similarities with traditional markets—risk-adjusted performance so far is broadly comparable, and the cross-section of returns can be summarized by a small set of factors. However, cryptocurrency also has its own distinct character: jumps are frequent and large, and blockchain information helps drive prices. This common set of stylized facts provides evidence that cryptocurrency is emerging as an investable asset class. Additionally, we discuss potential data quality issues and possible changes in future regulations and the cryptocurrency environment.
@@ 1 INTRODUCTION
This survey establishes a set of seven key empirical regularities of cryptocurrencies motivated by the view that cryptocurrency is an asset and thus can and should be analyzed from the asset-pricing point of view and, more specifically, from the cryptocurrency asset pricing view. Although the literature is relatively young—cryptocurrencies trace back to Nakamoto (2008)’s short paper on Bitcoin—it is now at the point of maturity where it is useful to establish a core of stylized facts in cryptocurrency markets that can be used for both empirical and theoretical work. This common set of facts also provides emerging evidence that cryptocurrency is becoming an investable asset class.
@@
When the paper of Nakamoto (2008) introduced a permissionless blockchain that provides public record-keeping without a central authority, financed and governed by a native token, the canonical literature on money had reached a degree of maturity. Permissionless ledgers can be viewed as technologies that supply public memory, with token systems financing and governing that record-keeping, and providing histories, communication, and enforcement mechanisms that are required for credit. With the rise of cryptocurrencies, these new ideas began to emerge primarily from practice more than theory and reopened classic monetary questions—record-keeping, commitment, enforcement—but now viewed from the point of view of blockchain economics.
@@
When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory. The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing. A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020; Liu, Tsyvinski & Wu 2022; Alvarez, Argente & Van Patten 2023; Sockin & Xiong 2023a; von Luckner, Reinhart & Rogoff 2023; Kogan et al. 2024).
@@
Many have viewed cryptocurrencies—and, not unjustifiably, many still view them—as hype, speculation, or fraud. When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system. [^1] See, Foley, Karlsen & Putniņš (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations. It is worth noting that several important financial innovations, such as the limited‑liability company and mortgage‑backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
@@
Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018; Leshno & Strack 2020; Biais et al. 2019; Kogan, Fanti & Viswanath 2022; Brunnermeier, James & Landau 2019; Benigno, Schilling & Uhlig 2022; Capponi, Jia & Wang 2025; Jermann & Xiang 2025), regulation (Cong & He 2019; Townsend 2020; Auer, Monnet & Shin 2025), digital asset pricing and valuation (Athey et al. 2016; Cong, Li & Wang 2021b; Biais et al. 2023a), transaction costs (Easley, O’Hara & Basu 2019; Cong, He & Li 2021a; Basu et al. 2023), central bank money (Auer et al. 2022; Whited, Wu & Xiao 2023; Schilling, Fernández-Villaverde & Uhlig 2024), securitization (Cong, Li & Wang 2022; Sockin & Xiong 2023a; Rogoff & You 2023; Goldstein, Gupta & Sverchkov 2024), safe assets (Ma, Zeng & Zhang 2025), international capital flow (von Luckner, Reinhart & Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al. 2024).
@@
We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of seven stylized facts—empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes. These stylized facts help evaluate and revisit both classic and new theories of money, while remaining conditional on the available data and institutional environment to date and therefore potentially evolving as market structure, regulation, and adoption change. [^2] For complementary surveys on blockchain architecture, decentralized finance and central bank digital currencies (CBDC), see Kose et al. (2022); Makarov et al. (2022); Kose et al. (2023); Biais et al. (2023b); Adrian & Mancini-Griffoli (2021); Infante et al. (2023).
@@
Seven Stylized Facts in Cryptocurrency Asset Pricing and Empirical Finance. Stylized fact 1 overviews the core asset-pricing characteristics of cryptocurrency markets—the risk-return trade-off and the diversification benefits. Stylized fact 2 discusses the structure of returns in cryptocurrency—factors, or “smart-beta”, which explain the cross-section of returns. Stylized fact 3 shows that the higher-order terms of a parsimonious four-factor model account for a large fraction of the cross-section of returns and provide a “glass box” alternative to the “black box” neural net models with interpretable economic factor structure. Stylized fact 4 shows the importance of the blockchain structure and network economy phenomena for cryptocurrency prices—on-chain user adoption influences cryptocurrency returns. Stylized fact 5 documents large and persistent market inefficiencies due to frictions that limit arbitrage, and discusses risks for arbitrageurs. Stylized fact 6 is about the cryptocurrency futures market and funding premia, and their evolution. Finally, stylized fact 7 is about the importance of regulation, disclosure, and transparency. In sum, these seven facts can be viewed as providing evidence that cryptocurrency, and more broadly, digital assets are coming of age as an emerging investable asset class.
@@ 2 DATA
Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf & Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025. The data on new addresses are from Liu, Tsyvinski & Wu (2021). Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025. We convert these quotes into U.S. dollars using spot FX rates from Bloomberg. Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.333Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S. dollar index, the Dow Jones US Real Estate Index, the U.S. constant maturity 10-year nominal and real bond yields from Bloomberg. The corresponding Bloomberg tickers are MXWO Index, LG30TRUU Index, LGTRTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
@@
Unless otherwise noted, we aggregate daily closing prices to construct weekly coin returns. [^4] CoinGecko daily prices are timestamped at 00:00 UTC (19:00 ET), whereas traditional asset prices from Bloomberg reflect end-of-day timestamps that vary by market and instrument (typically between 16:00 and 19:00 ET). Since our baseline analysis uses weekly close-to-close returns, any intraday timing mismatch is negligible. Crypto prices are quoted directly in USD, so no additional currency conversion is required. This review follows the dominant practice of using USD-denominated prices, which reflects the phenomenon that cryptocurrency trading is largely dominated by USD-linked quote currencies. We note that alternative numeraire or crypto-to-crypto return measures may reveal additional dimensions that are not captured by USD-based trades, which is a potentially important direction for future research. Each year is divided into 52 weeks: the first week contains the year’s first seven days, the next 50 weeks consist of seven days each, and the final week contains the remaining days of the year. For the construction of weekly returns, we exclude coins with a market capitalization below \$1 million at the formation date and winsorize returns at the 0.0025% level in both tails. We construct a cryptocurrency market return as the value-weighted return of all underlying coins. The cryptocurrency excess market return is constructed as the difference between the cryptocurrency market return and the risk-free rate measured as the 1-month T-bill rate at the weekly frequency, which we obtain from the Federal Reserve Bank of St. Louis.
An important caveat concerns data integrity. A strand of the empirical cryptocurrency literature argues that parts of the historical crypto data—especially in early years—may reflect manipulation or unreliable reporting, including alleged price manipulation on Mt. Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov & Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019). At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov & Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues. While these concerns cannot be fully eliminated, we mitigate them in several ways. First, our CoinGecko sample starts after the alleged Mt. Gox manipulation episode in 2013. Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov & Schoar (2020) and Cong et al. (2023b). Third, the CoinGecko prices we use are volume-weighted averages across a curated set of large spot exchanges and, per CoinGecko’s documentation, [^5] See, https://www.coingecko.com/en/methodology. are constructed using algorithms designed to detect and ignore outliers; as of the end of our sample these include the following exchanges: Binance, Coinbase, Kraken, Crypto.com, Bitfinex, Bitstamp, Gemini, and Huobi. Finally, our market-capitalization screens and data filters are designed to downweight or exclude small, thinly traded coins where data quality issues are more likely to be severe.
@@
Figure 1 summarizes the time-series evolution of cryptocurrency prices and market capitalization through September 2025. A one-dollar investment at the start of the sample grows to about \$160 in Bitcoin and \$80 in the overall market. There is substantial variation over time. Market capitalization reaches about \$4 trillion, with Bitcoin and Ethereum accounting for approximately 58% and 12% of the total at the end of the sample. [^6] Since the end of our sample, aggregate cryptocurrency market capitalization has declined; as of late January 2026 it is about \$2.8 trillion.
@@
3 THE SEVEN STYLIZED FACTS OF CRYPTOCURRENCY
We organize the main discussions of this review into seven key empirical regularities, or “stylized facts,” on cryptocurrency markets.
3.1 Stylized fact 1: High return, high volatility—normal Sharpe ratio so far, but rising comovement with equity
Stylized fact 1: The cryptocurrency market has an order of magnitude higher returns and volatility than equities, but a broadly similar Sharpe ratio; since 2020, it has also become substantially more correlated with equities.
@@ 3 THE SEVEN STYLIZED FACTS OF CRYPTOCURRENCY
Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable. Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.777These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk. Episodes such as exchange cybersecurity breaches or bankruptcies (e.g., Mt. Gox and FTX) can generate additional losses that are not reflected in spot price series.
Table 1, Panel A, reports summary statistics for the sample beginning in 2014 and ending in September 2025, as well as for the samples pre-2020 and post-2020. We compare returns on the aggregate coin market, a large-cap index of the top 100 cryptocurrencies (Coin 100), Bitcoin, and the aggregate global equity market. Across all three cryptocurrency indices, average weekly returns exceed 1% and volatilities are about an order of magnitude larger than for global equities; however, weekly Sharpe ratios are similar across cryptocurrencies and equities. Panel A also reports drawdowns and run-ups, which are substantially larger for cryptocurrencies than for equities. Importantly, drawdowns and return volatility for cryptocurrencies are smaller in the post-2020 sample than pre-2020, while the reverse holds for equities.
A complementary change concerns diversification. Table 1, Panel B, reports correlations between the coin market return and major asset classes. In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu & Tsyvinski (2021). In the post-2020 period, correlations rise sharply—most notably with global equities, increasing from 1% pre-2020 to 39% post-2020—and the coin market also becomes significantly correlated with commodities (27%), high-yield bonds (24%), government bonds (13%), and gold (18%). We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S. Treasury bonds with 10-year maturity. Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang & Lin (2023) and Jiang et al. (2025). Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization. The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market. [^8] Results are similar using a broad U.S. stock market index such as the S&P 500 (2% pre-2020; 37% post-2020), while correlations with U.S. government bonds (rather than global government bonds) remain insignificant in the post-2020 period.
These correlation patterns matter directly for portfolio choice. In mean–variance settings, an asset’s contribution to diversification depends on its covariance with the rest of the portfolio rather than its standalone volatility (Markowitz, 1952). Consistent with the portfolio results in Liu & Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set. We emphasize that these magnitudes are back-of-the-envelope and depend on strong assumptions about expected returns, correlations, and investability (liquidity, custody, and regulatory constraints), and should therefore be interpreted as illustrative upper bounds on diversification benefits rather than investment advice.
@@
Notes: This table summarizes the main properties of coin market return. Panel A reports weekly summary statistics of the coin market index and compares them with the value-weighted index of the top 100 coins by market capitalization (Coin 100), Bitcoin and global stock market returns. Reported measures include the mean, standard deviation, $t$-statistic, Sharpe ratio, skewness, kurtosis, maximum drawdown and run-up. Panel B presents pairwise correlation coefficients between coin market returns and returns on the global stock market, global high-yield bonds, global government bonds, a commodity index, gold, the dollar index, and the changes in inflation expectations. $t$-statistics are reported in parentheses. The full sample covers the period from January 7, 2014 to September 6, 2025. The pre-2020 sample is the sample before Jan 1, 2020, and the post-2020 sample is the sample since Jan 1, 2020.
3.2 Stylized fact 2: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value